OpenStudy (anonymous):

40c+15d distributive law of factor

OpenStudy (anonymous):

i factored it thats all i did and i got 10(4c+5d)

OpenStudy (anonymous):

@epoweritheta hun can u check to see if this is correct

OpenStudy (epoweritheta):

No @GotToLoveJenny , see we have ,40c+15d = 5*8c+5*3d (so we have 5 as a commoin factor) so , =5(8c+3d)

OpenStudy (anonymous):

okay thanks